OpenStudy (anonymous):

simplification of some logs, give me the fastest approach: log(10/10^m) and log(1/root4over10)

OpenStudy (anonymous):

 yeah this is the next step right? is it correct?

OpenStudy (anonymous):

But what do I do after?... it's just 1(10^1-m)

OpenStudy (alekos):

do you still need help?

OpenStudy (anonymous):

yes please sir

OpenStudy (anonymous):

@Auxuris :o

OpenStudy (anonymous):

oh oops do the same thing bring the 1-m to the left side of the log10 and then turn log10 into 1

OpenStudy (anonymous):

so it's supposed to be 1-m?

OpenStudy (anonymous):

but I still don't get how you move 1-m, I know it's that certain rule- but it just doesn't click.

OpenStudy (anonymous):

whenever your log or lg or ln has a power, you can bring it over to the other side

OpenStudy (anonymous):

but wouldn't it be 1-m(1)(10)?

OpenStudy (alekos):

it becomes (1-m)Log10

OpenStudy (anonymous):

what happened to the 10? from 10/10^m

OpenStudy (alekos):

10/10^m becomes 10^(1-m)

OpenStudy (alekos):

do you follow?

OpenStudy (anonymous):

log10(10^1-m)... so it's ... (1-m)(log10) but where did the 10 go?

OpenStudy (alekos):

Log(10/10^m) = Log(10^(1-m)) because 10/10^m is the same as 10^1/10^m which equals 10^(1-m) by the laws of exponents

OpenStudy (alekos):

can you see that now?

OpenStudy (anonymous):

not really.. ._. sorry, could you draw it? with arrows or something?

OpenStudy (alekos):

what don't you understand?

OpenStudy (anonymous):

where did the 10 go, because you need to multiply it out right?

OpenStudy (alekos):

no you don"t. 10/10^m = 10^(1-m) happy with that?

OpenStudy (anonymous):

then it's just log10(10^1-m) so it's 1(10^1-m)